# Physics-Regularized Multi-Modal Image Assimilation for Brain Tumor Localization

### v1

@@ Abstract
Physical models in the form of partial differential equations represent an important prior for many under-constrained problems. One example of such uses is tumor treatment planning, which heavily depends on an accurate estimate of the spatial distribution of tumor cells in a patient’s anatomy. Medical imaging scans can identify the bulk of the tumor, but they cannot reveal its full spatial distribution. Tumor cells at low concentrations remain undetectable, for example, in the most frequent type of primary brain tumors, glioblastoma. Deep-learning based approaches fail to estimate the complete tumor cell distribution due to a lack of reliable training data. Most existing works therefore rely on physics-based simulations to match observed tumors, providing anatomically and physiologically plausible estimations. However, these approaches struggle with complex and unknown initial conditions and are limited by overly rigid physical models. In this work, we present a novel method that allows balancing data-driven and physics-based cost functions. In particular, we propose a unique discretization scheme that allows quantifying the adherence of our learned spatiotemporal tumor and brain tissue distributions to their corresponding growth and elasticity equations. This quantification, serving as a regularization term rather than a hard constraint, enables greater flexibility and proficiency in assimilating patient data than existing models. We demonstrate improved coverage of tumor recurrence areas compared to existing techniques on real-world data from a cohort of patients. The method brings the potential to enhance clinical adoption of model-driven treatment planning for glioblastoma.
@@ 1 Introduction
The management of gliomas, notably glioblastomas, is highly challenging due to their infiltration beyond the tumor margins visible in medical scans such as Magnetic Resonance Imaging (MRI). This characteristic significantly complicates the accurate personalization of radiotherapy. Current clinical practices apply uniform distance margins around the visible tumor on MRI scans to target the full spatial distribution of tumor cells. However, these practices do not utilize our knowledge of infiltration patterns, tissue-specific characteristics, and anatomical barriers, thereby failing to capture the complex dynamics of individual tumor spread. Effectively tailoring radiotherapy to an individual patient’s tumor’s spread is an unmet clinical need in neurooncology [1].
To address this challenge, two main computational approaches have emerged:
@@
These models use Partial Differential Equations (PDEs) to constrain the solution space, ensuring that output tumor cell distributions adhere strictly to predefined physical laws [2, 3, 4, 5]. However, since PDEs are only approximations of the underlying stochastic processes in complex biological environments, they can overconstrain the system. While such models can be informative and structured, they cannot capture the complex behaviors intrinsic to biological processes like tumor growth. Consequently, the data cannot be fully assimilated due to these restrictions, leading to a limited utility of the physical models in representing empirical data.
@@
These models [6, 7] offer great flexibility by leveraging access to imaging datasets. However, they often lack rigorous quality control and fail to integrate crucial biological insights, such as tissue-specific infiltration patterns and brain topology. This makes them less reliable for clinical application, especially when there is insufficient data to infer them purely from imaging information.
Combining physical insights on spatiotemporal distributions with increased model flexibility could be highly beneficial. One approach to achieve this is by testing the physics residual in selected locations, as demonstrated in Physics-Informed Neural Networks (PINNs) [8, 9, 10, 11, 12]. Another way involves projecting the learned solution into a latent space that adheres to physics constraints [13, 14]. A different framework, Optimizing a Discrete Loss (ODIL) [15], constructs a residual of PDEs using grid-based discretization. Unlike methods that optimize neural network parameters, ODIL directly optimizes the discretized unknown fields
Both ODIL [16] and PINNs [17] have recently been employed to model growth equations conditioned on MRI and FET-PET imaging of patients. ODIL advances state-of-the-art radiotherapy planning by relaxing the constraints of the growth equation, thereby enhancing the capture of tumor recurrence. A similar trend is observed in computer vision’s novel view synthesis, where neural radiance field MLPs [18] are being replaced by discrete representations [19, 20, 21].
However, these approaches do not account for brain tissues as part of the dynamics and do not utilize the information available in the imaging data regarding visible anatomical deformations. Incorporating tissue elasticity is challenging because the exact initial anatomy is unknown, and the dynamics themselves are computationally challenging due to the significantly increased number of unknown fields. Thus, no hybrid approach free from hard PDE constraints has yet successfully connected biological processes, such as tumor growth, with models of its biological environment, such as the tissue elasticity models. Very few works have been done [22, 23] using PINNs for biomechanical modeling only on synthetic scenarios, demonstrating that the displacement field of materials can be learned from observations.
@@
  • We construct discrete physics-based residuals for both the tumor growth and the biomechanical processes of dynamic tissues from the Lagrangian perspective of particles that carry tissue intensities. Eulerian projections on static grids are used to condition the learnable tumor cell and tissue distributions on available visible observations. Unlike data-driven models, our method remains robust due to physics regularization while being more flexible at adapting data than hard-constrained numerical simulations.
@@
  • We validate our method  [^1] GitHub: https://github.com/m1balcerak/PhysRegTumor on the downstream task of radiotherapy planning using the largest publicly available dataset known to us, achieving new state-of-the-art performance in capturing tumor recurrence.
Figure 1: Method overview. (a) 3D MRI and PET scans of a glioblastoma patient (b) Preprocessed input includes brain tissue maps, tumor segmentation, and a metabolic map from FET-PET. (c) Tumor cell distribution and brain anatomy inferred using a loss function based on assumptions about physical processes, initial conditions. (d) Outcomes: initial healthy anatomy, spatial tumor cell distribution, and system identification parameters.
In Figure 1, we visualize an overview of our method, showing how we condition the learnable full spatial tumor cell distribution on the patient data and regularize it using soft assumptions about the physics of tumor growth, tissue elasticity, and initial conditions for both the anatomy and the tumor. In addition to multiparametric MRI scans (T1Gd, T1, FLAIR), we incorporate Fluoroethyl-L-tyrosine PET (FET-PET), a specialized form of Positron Emission Tomography (PET), as an additional imaging modality that captures the metabolic activity of tumor cells.
@@ 2.1 Physical Model
The brain tissue vector $\boldsymbol{m}$ is governed by
@@
The tumor induced stresses lead to a displacement field $\boldsymbol{u}$, such that the velocity is given by $\boldsymbol{v}=\partial_{t}\boldsymbol{u}$. We assume quasi-static mechanical equilibrium due to the slow tumor growth rate relative to tissue mechanical responses, leading to
@@
where $\gamma$ regulates the impact of the tumor on the displacement and is a learnable patient-specific parameter. We apply the Neo-Hookean model [25] to represent hyperelastic behavior in biological tissues, where the stress $\boldsymbol{\sigma}$ is related to the deformation gradient $\mathbf{F}_{ij}=\delta_{ij}+\frac{\partial u_{i}}{\partial x_{j}}$ with the Lamé parameters $\lambda$ and $\mu$ (material properties in Appendix E), averaged across different tissue types to accommodate the heterogeneity:
@@ 2.2 Discrete Physics Residuals
We use two different approaches to discretize the elasticity equations and the reaction-diffusion-advection equations. For the former, we use a time-dependent grid in space, described by the $N_{x}$ points $(\boldsymbol{p}^{n}_{j})_{1\leq j\leq N_{X}}$ at time $n\Delta t$, $0\leq n\leq N_{t}$, which initially form a uniform grid on $\Omega$. For the latter, we further partition the domain $\Omega$ at the $n$th point in time into $N_{x}$ cells $\Omega_{i}^{n}$, $1\leq i\leq N_{x}$, such that $\Omega=\bigcup_{i=1}^{N_{x}}\Omega_{i}^{n}$. Note that the cell midpoints $\boldsymbol{x}_{j}$ directly depend on the particle positions $\boldsymbol{p}_{j}$ due to the tissue deformation and that the shape of the cells $\Omega_{i}$ is given by a Voronoi tesselation of $\boldsymbol{x}_{j}$.
The displacement field $\boldsymbol{u}^{n}_{j}$ is computed based on the initial ($t=0$) and current ($t=n\Delta t$) positions of particles $\boldsymbol{p}^{n}$ within the tissue by $\boldsymbol{u}^{n}=\boldsymbol{p}^{n}-\boldsymbol{p}^{0}$. Using central finite differences on the initial uniform reference grid, we can consecutively approximate $\nabla\boldsymbol{u}_{j}^{n}$, $\boldsymbol{F}^{n}_{j}$, $\nabla c^{n}_{j}$, $\nabla\cdot\boldsymbol{\sigma}^{n}_{j}$, and define the tissue residual
@@
We use an implicit Euler scheme combined with a cell-centered finite-volume discretization to discretize the tumor equation. We denote the boundaries between two cells $1\leq i,j\leq N_{x}$ by $\Gamma_{ij}^{n}=\Gamma_{ji}^{n}=\partial\Omega_{i}^{n}\cap\partial\Omega_{j}^{n}$. Further, the index set of all neighboring cells of $\Omega^{n}_{i}$ is given by $N_{i}=\{j:\partial\Omega_{j}\cap\partial\Omega_{i}\neq\varnothing\land j\neq i\}$ and assumed to stay independent of $n$, due to the extend of our deformations (see limitations in Section 7). For a function $c^{n}_{i}$ at time $n\Delta t$ at the $i$th cell, we discretize the diffusion $\mathcal{D}$ and reaction $\mathcal{S}$ operators by
$\mathcal{D}[c^{n}_{i},D^{n}_{i}]=\sum_{j\in N_{i}}|\Gamma_{ij}|D_{ij}^{n}\frac{c_{j}^{n}-c_{i}^{n}}{\lVert\boldsymbol{x}_{j}^{n}-\boldsymbol{x}_{i}^{n}\rVert},\textmd{ and }\mathcal{S}[u^{n}_{i}]=|\Omega_{i}^{n}|\rho c_{i}^{n}(1-c_{i}^{n}),$ (6)
where $D_{ij}^{n}$ is the harmonic mean of $D^{n}$ at the cell boundary, $|\Gamma_{ij}^{n}|$ denotes the interface area, and $|\Omega_{i}^{n}|$ denotes the cell volume. We thus obtain the tumor loss by
$L_{\text{Tumor Growth PDE}}=\sum_{n=0}^{N_{t}}\sum_{j=0}^{N_{x}}\left(\frac{|\Omega^{n}_{i}|}{\Delta t}(c^{n}_{i}-c^{n-1}_{i})-\mathcal{D}[c^{n},D^{n}]-\mathcal{S}[c^{n}]\right)^{2}.$ (7)
Figure 2: Learning process overview: optimization of spatiotemporal distributions of tumor cells and tissues. (a) Initial condition penalties enforce symmetric healthy anatomy, with the initial tumor distribution at $t=0$ as a Gaussian. (b) Physics penalties regularize dynamics between $t=0$ and $t_{\text{Final}}$. The first row shows gray matter contours with particle positions; the second row shows white matter contours with learned tumor concentrations. (c) Agreement of tumor distribution with anatomical tissues, visible segmentations, and metabolic map after transforming the final tumor distribution through the imaging function.
@@ 2.3 Particle-Grid Projections
$\mathcal{F}_{p}^{P}\approx\sum_{i}w_{ip}\mathcal{F}_{i}^{G},\quad\mathcal{F}_{i}^{G}\approx\frac{\sum_{p}w_{ip}\mathcal{F}_{p}^{P}}{\sum_{p}w_{ip}}$ (8)
where $i$ indexes grid nodes, $p$ indexes particles, and $w_{ip}$ represents the weight of the trilinear shape function defined on node $i$ and evaluated at the location of particle $p$. .
@@ 3 Initial Assumptions
To model the initial conditions of the tumor and brain tissues, we incorporate specific assumptions into our loss functions. These assumptions help to align the model with expected biological and anatomical conditions without the tumor and at the time of its origin.
@@ 3.1 Initial Tumor Distribution
We assume that the initial tumor distribution follows a Gaussian profile. This is represented by penalizing the difference between $\mathcal{F}^{G}(u)^{0}$ and a Gaussian function centered at $x_{0}\in\Omega$. The center of this Gaussian profile, $x_{0}$, is part of the learnable parameters $\theta_{\text{Initial}}=\{x_{0},\boldsymbol{m}^{0}\}$. The initial condition loss $L_{\text{Tumor starts from a Gaussian}}$ is as defined as:
$L_{\text{Tumor starts from a Gaussian}}=\sum_{i=0}^{N_{x}}\left(\mathcal{F}^{G}(c)^{0}_{i}-D_{1}\exp\left(-\frac{(x_{i}-x_{0})^{2}}{D_{2}}\right)\right)^{2}$ (9)
where $D_{1}=0.5$ and $D_{2}=0.02$ are constants.

### v3

@@ Abstract
Physical models in the form of partial differential equations serve as important priors for many under-constrained problems. One such application is tumor treatment planning, which relies on accurately estimating the spatial distribution of tumor cells within a patient’s anatomy. While medical imaging can detect the bulk of a tumor, it cannot capture the full extent of its spread, as low-concentration tumor cells often remain undetectable, particularly in glioblastoma, the most common primary brain tumor. Machine learning approaches struggle to estimate the complete tumor cell distribution due to a lack of appropriate training data. Consequently, most existing methods rely on physics-based simulations to generate anatomically and physiologically plausible estimations. However, these approaches face challenges with complex and unknown initial conditions and are constrained by overly rigid physical models. In this work, we introduce a novel method that integrates data-driven and physics-based cost functions, akin to Physics-Informed Neural Networks (PINNs). However, our approach parametrizes the solution directly on a dynamic discrete mesh, allowing for the effective modeling of complex biomechanical behaviors. Specifically, we propose a unique discretization scheme that quantifies how well the learned spatiotemporal distributions of tumor and brain tissues adhere to their respective growth and elasticity equations. This quantification acts as a regularization term, offering greater flexibility and improved integration of patient data compared to existing models. We demonstrate enhanced coverage of tumor recurrence areas using real-world data from a patient cohort, highlighting the potential of our method to improve model-driven treatment planning for glioblastoma in clinical practice.
@@ 1 Introduction
The management of gliomas, particularly glioblastomas, is highly challenging due to their infiltration beyond the tumor margins detectable in medical imaging. This infiltrative behavior complicates the precise personalization of radiotherapy. Current clinical practice applies uniform radiation to a 1.5 cm margin around the active tumor core visible in Magnetic Resonance Imaging (MRI) scans. However, this approach overlooks critical factors such as heterogeneous infiltration patterns, varying brain tissues, and anatomical barriers, failing to account for the highly individualized dynamics of tumor spread in each patient. Personalizing radiotherapy to account for this spread remains an unmet clinical need in neuro-oncology [1].
The computational modeling of tumor growth has become a promising solution to this challenge [2, 3, 4, 5], by formalizing the tumor growth process through mathematical models of varying complexity [6]. However, personalizing these models to the limited clinical observational (image) data of a patient remains a formidable challenge. To address it, two main computational approaches have emerged:
@@
These models use Partial Differential Equations (PDEs) to constrain the solution space, ensuring that output tumor cell distributions adhere strictly to predefined physical laws [7, 8, 9, 10, 11].
However, since PDEs are approximations of the underlying stochastic processes in complex biological environments, they may overconstrain the system. While these models offer structure and insight, they fail to capture the intricate behaviors inherent to biological processes like tumor growth. Additionally, linking image observations to state variables requires careful consideration. As a result, the full integration of data is hindered by these limitations, reducing the effectiveness of physical models in representing empirical data.
@@
These models [12, 13] offer great flexibility by leveraging access to imaging datasets. They can directly be learned from – and applied to – the patients’ diagnostic scans. However, they often lack rigorous quality control and fail to integrate crucial biological insights, such as tissue-specific infiltration patterns and brain topology. This makes them less reliable for clinical application, especially when there is insufficient data to infer them purely from imaging information.
In Search of a Hybrid Approach:
Combining physical insights on spatiotemporal distributions of state variables with increased model flexibility could be highly beneficial. One approach to achieve this is by testing the physics residual in selected locations, as demonstrated in Physics-Informed Neural Networks (PINNs) [14, 15, 16, 17, 18]. Another way involves projecting the learned solution into a latent space that adheres to physics constraints [19, 20]. A different framework, Optimizing a Discrete Loss (ODIL) [21], constructs a residual of PDEs using grid-based discretization. Unlike methods that optimize the weights of Multilayer Perceptrons (MLPs) that implicitly parametrize the unknown fields in the PINNs method, ODIL directly optimizes the discretized unknown fields, which can be faster and more stable due to the local (rather than global, as in MLPs) dependencies of the solution on the learnable weights.
Both ODIL [22] and PINNs [23] have recently been employed to model growth equations conditioned on MRI and metabolic imaging of patients. ODIL advances state-of-the-art radiotherapy planning by relaxing the constraints of the growth equation, thereby enhancing the capture of tumor recurrence. A similar trend is observed in computer vision’s novel view synthesis, where neural radiance field MLPs [24] are being replaced by discrete representations [25, 26, 27].
However, none of these hybrid approaches have modeled brain-tumor interactions dealing with constraints imposed by local structural anatomy, such as the relative composition of brain tissues and their microstructure, or by the impact of gross anatomical deformations visible in structural scans of large tumors. While biomechanical model extensions can address the latter, incorporating tissue elasticity remains challenging due to the unknown initial anatomy and the computational complexity of the tumor growth dynamics. So far, no hybrid approach free from hard PDE constraints has successfully linked biological processes, such as tumor growth, with models of the biological environment, like tissue elasticity. Only a few studies have used PINNs for biomechanical modeling in synthetic scenarios, demonstrating that the displacement field of materials can be learned from observations [28, 29].
@@
  • We construct discrete physics-based residuals as measures of differences between learned spatiotemporal distributions and those predicted by discretized physical equations governing tumor growth and tissue biomechanics. Using a Lagrangian perspective where particles carry tissue intensities, we project these dynamics onto static Eulerian grids to condition the learnable tumor cell and tissue distributions on available observations. Unlike data-driven models, our hybrid method remains robust due to physics regularization while being more flexible at adapting data than hard-constrained numerical simulations.
@@
  • We validate our method [^1] Code is available at https://github.com/m1balcerak/PhysRegTumor on the downstream task of radiotherapy planning using the largest publicly available dataset known to us, achieving new state-of-the-art performance in capturing tumor recurrence.
Figure 1: Method overview: (a) 3D MRI and PET scans of a glioblastoma patient (b) Preprocessed input includes brain tissue maps, tumor segmentation, and a metabolic map from FET-PET. (c) Tumor cell distribution and brain anatomy inferred using a loss function based on assumptions about physical processes, initial conditions. (d) Outcomes: initial healthy anatomy, spatial tumor cell distribution, and system identification parameters.
Figure 1 provides an overview of our method, showing how we condition the learnable full spatial tumor cell distribution on the patient data and regularize it using soft assumptions about the physics of tumor growth, tissue elasticity, and initial conditions for both the anatomy and the tumor. In addition to multiparametric MRI scans (T1Gd, T1, FLAIR), we incorporate Fluoroethyl-L-tyrosine PET (FET-PET), a specialized form of Positron Emission Tomography (PET), as an additional imaging modality that captures the metabolic activity of tumor cells.
@@ 2.1 Physical Model
The brain tissue vector $\boldsymbol{m}$ is governed by:
@@
The tumor induced stresses lead to a displacement field $\boldsymbol{u}$, such that the velocity is given by $\boldsymbol{v}=\partial_{t}\boldsymbol{u}$. We assume quasi-static mechanical equilibrium due to the slow tumor growth rate relative to tissue mechanical responses, leading to:
@@
where $\gamma$ regulates the impact of the tumor on the displacement and is a learnable patient-specific parameter. We apply the Neo-Hookean model [31] to represent hyperelastic behavior in biological tissues, where the stress $\boldsymbol{\sigma}$ is related to the deformation gradient $\mathbf{F}_{ij}=\delta_{ij}+\frac{\partial u_{i}}{\partial x_{j}}$ with the Lamé parameters $\lambda$ and $\mu$ (material properties in Appendix G), averaged across different tissue types to accommodate the heterogeneity:
@@ 2.2 Discrete Physics Residuals
We use two different approaches to discretize the elasticity equations and the reaction-diffusion-advection equations. For the former, we use a time-dependent grid in space, described by the $N_{x}$ points $(\boldsymbol{p}^{n}_{j})_{1\leq j\leq N_{X}}$ at time $n\Delta t$, $0\leq n\leq N_{t}$, which initially form a uniform grid on $\Omega$. For the latter, we further partition the domain $\Omega$ at the $n$th point in time into $N_{x}$ cells $\Omega_{i}^{n}$, $1\leq i\leq N_{x}$, such that $\Omega=\bigcup_{i=1}^{N_{x}}\Omega_{i}^{n}$. Note that the nodes defining the cell boundaries are carried by the particles $\boldsymbol{p}^{n}_{j}$, whose movement, governed by tissue elasticity, gradually deforms the cells $\Omega^{n}_{i}$ from their initial rectangular shape.
The displacement field $\boldsymbol{u}^{n}$ is computed based on the initial ($t=0$) and current ($t=n\Delta t$) positions of particles $\boldsymbol{p}^{n}$ within the tissue by $\boldsymbol{u}^{n}=\boldsymbol{p}^{n}-\boldsymbol{p}^{0}$. Using central finite differences on the initial uniform reference grid, we can consecutively approximate $\nabla\boldsymbol{u}_{j}^{n}$, $\boldsymbol{F}^{n}_{j}$, $\nabla c^{n}_{j}$, $\nabla\cdot\boldsymbol{\sigma}^{n}_{j}$, and define the tissue residual:
@@
We use an implicit Euler scheme combined with a cell-centered finite-volume discretization to discretize the tumor equation. We denote the boundaries between two cells $1\leq i,j\leq N_{x}$ by $\Gamma_{ij}^{n}=\Gamma_{ji}^{n}=\partial\Omega_{i}^{n}\cap\partial\Omega_{j}^{n}$. Further, the index set of all neighboring cells of $\Omega^{n}_{i}$ is given by $N_{i}=\{j:\partial\Omega_{j}\cap\partial\Omega_{i}\neq\varnothing\land j\neq i\}$ and assumed to stay independent of $n$, due to the extend of our deformations (see limitations in Section 7). For a function $c^{n}_{i}$ at time $n\Delta t$ at the $i$th cell, we discretize the diffusion $\mathcal{D}$ and reaction $\mathcal{S}$ operators by:
$\mathcal{D}[c^{n}_{i},D^{n}_{i}]=\sum_{j\in N_{i}}|\Gamma_{ij}|D_{ij}^{n}\frac{c_{j}^{n}-c_{i}^{n}}{\lVert\boldsymbol{x}_{j}^{n}-\boldsymbol{x}_{i}^{n}\rVert},\textmd{ and }\mathcal{S}[c^{n}_{i}]=|\Omega_{i}^{n}|\rho c_{i}^{n}(1-c_{i}^{n}),$ (6)
where $D_{ij}^{n}$ is the harmonic mean of $D_{i}^{n}$ at the cell boundary, $|\Gamma_{ij}^{n}|$ denotes the interface area, and $|\Omega_{i}^{n}|$ denotes the cell volume. We thus obtain the tumor loss by:
$L_{\text{Tumor Growth PDE}}=\sum_{n=1}^{N_{t}}\sum_{i=0}^{N_{x}}\left(\frac{|\Omega^{n}_{i}|}{\Delta t}(c^{n}_{i}-c^{n-1}_{i})-\mathcal{D}[c^{n}_{i},D^{n}_{i}]-\mathcal{S}[c^{n}_{i}]\right)^{2}.$ (7)
Figure 2: Learning process overview: optimization of spatiotemporal distributions of tumor cells and tissues. (a) Initial condition penalties enforce symmetric healthy anatomy, with the initial tumor distribution at $t=0$ as a small Gaussian blob. (b) Physics penalties regularize dynamics between the initial and the final time. The first row shows gray matter contours with particle positions; the second row shows white matter contours with learned tumor concentrations. (c) Agreement of tumor distribution with anatomical tissues, visible tumor segmentations, and metabolic map after transforming the final tumor distribution through the imaging function.
@@ 2.3 Particle-Grid Projections
$\mathcal{F}_{j}^{P}\approx\sum_{i}w_{ij}\mathcal{F}_{i}^{G},\quad\mathcal{F}_{i}^{G}\approx\frac{\sum_{p}w_{ij}\mathcal{F}_{j}^{P}}{\sum_{j}w_{ij}},$ (8)
where $i$ indexes grid nodes, $j$ indexes particles, and $w_{ij}$ represents the weight of the trilinear shape function defined on node $i$ and evaluated at the location of particle $j$.
@@ 3 Initial Assumptions
We model the initial conditions of the tumor and brain tissues by incorporating assumptions into our loss functions, ensuring the model aligns with the expected biological and anatomical state of the brain before and at the onset of tumor growth.
@@ 3.1 Initial Tumor Distribution
We assume that the initial tumor distribution can be represented by a small Gaussian blob. To enforce this assumption, we penalize the difference between $\mathcal{F}^{G}(u)^{0}$ and a Gaussian function centered at $x_{0}\in\Omega$. The center of this Gaussian blob, $x_{0}$, is part of the learnable parameters $\theta_{\text{Initial}}=\{x_{0},\boldsymbol{m}^{0}\}$. The initial condition loss for the tumor cells is defined as:
$L_{\text{Tumor starts from a Gaussian}}=\sum_{i=0}^{N_{x}}\left(\mathcal{F}^{G}(c)^{0}_{i}-D_{1}\exp\left(-\frac{(x_{i}-x_{0})^{2}}{D_{2}}\right)\right)^{2},$ (9)
where $D_{1}=0.5$, $D_{2}=0.02$ are constants and $x_{i}$ is the position of the center of $\Omega^{0}_{i}$.